# Gait in Eight: Efficient On-Robot Learning for Omnidirectional Quadruped Locomotion

### v1

@@ I INTRODUCTION
Legged robot locomotion has long been an important research area in robotics, as achieving robust, agile, and adaptable gaits in unstructured environments is a challenging yet essential capability for many real-world applications. Traditional model-based controllers [1, 2], while effective in well-structured settings, often struggle to handle the inherent uncertainties and dynamic variations encountered in real-world terrains. In contrast, Deep Reinforcement Learning (DRL) offers a promising paradigm that allows robots to autonomously acquire locomotion skills directly through interaction with the environment. In recent years, DRL has shown remarkable success in learning complex and agile locomotion skills for many different legged robots [3], such as high-speed running [4, 5], jumping and climbing in parkour-like courses [6, 7], navigating through challenging terrain [8, 9], and performing handstands and backflips [7, 10]. However, these works rely on scaling up on-policy DRL algorithms, mainly Proximal Policy Optimization (PPO) [11], through thousands of parallel simulated environments with GPU-based physics engines [12]. While a plethora of domain randomization is necessary to zero-shot transfer policies trained in simulation to the real world [13, 14], this creates a significant embodiment gap between the widely randomized and approximated dynamics of the simulated robot and the specific, nuanced dynamics of the real robot. On-robot learning promises to bridge the embodiment gap by learning directly on the real system. This enables the DRL agent to be aware of its physical embodiment and the specific hardware constraints. It can continuously adapt to changes, such as wear and tear, battery depletion, hardware modifications, or environmental changes. Recent advances in off-policy DRL have made first steps toward this paradigm by improving the learning efficiency enough to enable training quadruped locomotion in real-time directly on the robot [15, 16, 17]. However, these works must simplify the learning task to plain forward locomotion with a fixed target velocity. They only achieve crawling gaits and rely on powerful laptops with dedicated GPUs to run the training process fast enough to be feasible. Our goal is to lift computational constraints by improving the learning efficiency and update speed with the recently proposed CrossQ DRL algorithm [18] and extend the learning task to omnidirectional locomotion, allowing any desired velocity in the -plane.
@@ II-C Quadruped platform
III EFFICIENT FOR ON-ROBOT LEARNING FOR LEGGED LOCOMOTION
@@ V CONCLUSION
In this work, we presented a framework for efficiently learning quadruped locomotion directly on the HoneyBadger quadruped robot using the CrossQ algorithm. Our approach leverages CrossQ’s sample efficiency and minimal compute overhead to achieve maximum velocities of up to in just 8 minutes of raw training. We combine CrossQ with two control architectures: a JTP scheme for agile, high-speed gaits and a CPG scheme for stable, natural gaits. Lastly, we extended on-robot locomotion learning to omnidirectional locomotion with different target velocities in the -plane. Our real-world experiments in indoor and outdoor environments showed the practicality of our framework and the robustness of the learned policies to terrain variations and sensor noise. Future work will focus on improving the linear velocity estimation, exploring visual observations, and fine-tuning powerful pre-trained policies from simulation to adapt to new environments and the specific robot embodiment that is changing through wear and tear or hardware modifications.

### v2

@@ I INTRODUCTION
Legged robot locomotion has long been an important research area in robotics, as achieving robust, agile, and adaptable gaits in unstructured environments is a challenging yet essential capability for many real-world applications. Traditional model-based controllers [1, 2], while effective in well-structured settings, often struggle to handle the inherent uncertainties and dynamic variations encountered in real-world terrains. In contrast, Deep Reinforcement Learning (DRL) offers a promising paradigm that allows robots to autonomously acquire locomotion skills directly through interaction with the environment. In recent years, DRL has shown remarkable success in learning complex and agile locomotion skills for many different legged robots [3, 4], such as high-speed running [5, 6], jumping and climbing in parkour-like courses [7, 8], navigating through challenging terrain [9, 10], and performing handstands and backflips [8, 11]. However, these works rely on scaling up on-policy DRL algorithms, mainly Proximal Policy Optimization (PPO) [12], through thousands of parallel simulated environments with GPU-based physics engines [13]. While a plethora of domain randomization is necessary to zero-shot transfer policies trained in simulation to the real world [14, 15], this creates a significant embodiment gap between the widely randomized and approximated dynamics of the simulated robot and the specific, nuanced dynamics of the real robot. On-robot learning promises to bridge the embodiment gap by learning directly on the real system. This enables the DRL agent to be aware of its physical embodiment and the specific hardware constraints. It can continuously adapt to changes, such as wear and tear, battery depletion, hardware modifications, or environmental changes. Recent advances in off-policy DRL have made first steps toward this paradigm by improving the learning efficiency enough to enable training quadruped locomotion in real-time directly on the robot [16, 17, 18]. However, these works must simplify the learning task to plain forward locomotion with a fixed target velocity. They only achieve crawling gaits and rely on powerful laptops with dedicated GPUs to run the training process fast enough to be feasible. Our goal is to lift computational constraints by improving the learning efficiency and update speed with the recently proposed CrossQ DRL algorithm [19] and extend the learning task to omnidirectional locomotion, allowing any desired velocity in the -plane.
@@ II-C Quadruped platform
III EFFICIENT ON-ROBOT LEARNING FOR LEGGED LOCOMOTION
@@ V CONCLUSION
In this work, we presented a framework for efficiently learning quadruped locomotion directly on the HoneyBadger quadruped robot using the CrossQ algorithm. Our approach leverages CrossQ’s sample efficiency and minimal compute overhead to achieve maximum velocities of up to in just 8 minutes of raw training. We combine CrossQ with two control architectures: a JTP scheme for agile, high-speed gaits and a CPG scheme for stable, natural gaits. Lastly, we extended on-robot locomotion learning to omnidirectional locomotion with different target velocities in the -plane. Our real-world experiments in indoor and outdoor environments showed the practicality of our framework and the robustness of the learned policies to terrain variations and sensor noise. Future work will focus on improving the linear velocity estimation with visual-inertial odometry [39], exploring visual observations, algorithmic improvements of the base algorithm [40] and fine-tuning powerful pre-trained policies from simulation to adapt to new environments and the specific robot embodiment that is changing through wear and tear or hardware modifications.